OpenStudy (anonymous):

cashews cost $5 per pound and the peanuts cost 1.50 per pound. how many pounds of cashews should be mixed with peanuts to get a mixture of 32 pounds that will cost $4 per pound? Round to the nearest tenth.

OpenStudy (anonymous):

x+y=32......................(1) 5x+1.5y=32*4.............(2) 5x+1.5y=128 y=32-x......................(1) Substitude (1) in (2): 5x+1.5(32-x)=128 5x+48-1.5x=128 3.5x=80 x=22.9 pounds of peanuts y=32-22.9=9.1 pounds of cashews

OpenStudy (anonymous):

that was fast thank you